# Adaptive Online Bayesian Estimation of Frequency Distributions with Local Differential Privacy

### v1

@@ 3 Problem definition and general framework
Suppose we are interested in a discrete probability distribution $\mathcal{P}$ of a certain form of sensitive categorical information $X\in\mathcal{X}:=\{1,\ldots,K\}$ of individuals in a population. Hence, $\mathcal{P}$ is a categorical distribution $\text{Cat}(\theta^{\ast})$ with a probability vector
@@
$\Delta:=\left\{\theta\in\mathbb{R}^{K}:\sum_{k=1}^{K}\theta_{k}=1\text{ and }\theta_{k}\geq 0\text{ for }k\in\{1,\ldots,K\}\right\}.$
We assume a setting where the individuals’ sensitive data are collected privately and sequentially in time. The privatization is performed via a randomized algorithm that, upon taking a category index in $\{1,\ldots,K\}$ as an input, returns a random category index in $\{1,\ldots,K\}$ such that the whole data collection process is $\epsilon$-LDP. (See Definition 1 above.) Let $X_{t}$ and $Y_{t}$ be the private information and randomized responses of individual $t$, respectively. According to Definition 1 for LDP, the following inequality must be satisfied for all triples $(x,x^{\prime},y)\in\mathcal{X}^{3}$ for the randomized mechanism to be $\epsilon$-LDP.
@@
This paper concerns online Bayesian estimation of $\theta$ while adapting the randomized mechanism to improve the estimation utility continually. We propose a general framework called Adaptive Online Bayesian Estimation with LDP (AdOBEst-LDP), in which the randomized mechanism at time $t+1$ is adapted to the data collected until time $t$. AdOBEst-LDP is outlined in Algorithm 1.
@@ 5      Step 3: Draw a sample (approximately) from the posterior distribution
Algorithm 1 is fairly general and it does not describe how to choose the $\epsilon$-LDP mechanism $\mathcal{M}_{t}$ at time $t$, nor does it provide the details of the posterior sampling. However, it is still worth making some critical observations about the nature of the algorithm. Firstly, note that the selection of the $\epsilon$-LDP mechanism in Step 1 relies on the posterior sample $\Theta_{t}$, which serves as an estimator of the true parameter $\theta^{\ast}$. As we shall see in Section 4, at Step 1 the ‘best’ $\epsilon$-LDP mechanism is chosen from a set of candidate LDP mechanisms according to a utility function. This step is relevant only when $\Theta_{t}$ is a reliable estimator of $\theta^{\ast}$. In other words, Step 1 ‘exploits’ the estimator $\Theta_{t}$. Moreover, the random nature of posterior sampling prevents having too much confidence in the current estimator $\Theta_{t}$ and enables a certain degree of ‘exploration’. In conclusion, Algorithm 1 utilizes an ‘exploration-exploitation’ approach reminiscent of reinforcement learning. In particular, posterior sampling in Step 3 suggests a strong parallelism between AdOBEst-LDP and the well-known exploration-exploitation approach called Thompson sampling (Russo et al., , 2018).
@@ 4 Constructing informative randomized response mechanisms
In this section, we describe Step 1 of AdOBEst-LDP in Algorithm 1 where the $\epsilon$-LDP mechanism $\mathcal{M}_{t}$ is selected at time $t$ based on the posterior sample $\Theta_{t}$. For ease of exposition, we will drop the time index $t$ throughout the section and let $\Theta_{t}=\theta$.
Recall that an $\epsilon$-LDP randomized mechanism is associated with a conditional probability distribution that satisfies (1). An $\epsilon$-LDP mechanism is not unique. One such mechanism, which we call Standard Randomized Response (SRR), returns $Y=X$ with probability $e^{\epsilon}/(e^{\epsilon}+K-1)$, otherwise returns a uniformly random number on $\{1,\ldots,K\}/\{X\}$. The transition law corresponding to the SRR from $X$ to $Y$ is
$\mathbb{P}_{\text{SRR}}(Y=y|X=x)=\frac{e^{\epsilon}}{e^{\epsilon}+K-1}\mathbb{I}(x=y)+\frac{1}{e^{\epsilon}+K-1}\mathbb{I}(x\neq y).$
For subsequent use, we define $\texttt{SRR}(\epsilon,\mathcal{X})$ to be the SRR mechanism that operates on a sample space $\mathcal{X}$ with DP parameter $\epsilon$, and $\texttt{SRR}(x|K,\epsilon)$ to be its random output when it takes $X=x$ as its input.
We aim to develop an alternative randomized mechanism such that the generated randomized response $Y$ is more informative about $\theta^{\ast}$ than the one generated by SRR. The main idea is as follows. Supposing that the posterior sample $\Theta_{t}=\theta$ is an accurate estimate of $\theta^{\ast}$, it is reasonable to aim for the ‘best’ $\epsilon$-LDP mechanism (among a range of candidates) which would maximize the (estimation) utility of $Y$ if the true parameter were $\theta^{\ast}=\theta$. We follow this main idea to develop the proposed $\epsilon$-LDP mechanism.
@@ 4.1 The randomly restricted randomized response (RRRR) mechanism
Given $\Theta_{t}=\theta\in\Delta$, an informative randomized response mechanism can be constructed by considering a high-probability set $S\subset\{1,\ldots,K\}$ for $X$, and a low-probability set $S^{c}=\{1,\ldots,K\}/S$ for $X$. Let $\sigma_{\theta}=\sigma_{\theta}(1),\ldots,\sigma_{\theta}(K)$ be a permutation vector for $\theta$ so that $\theta_{\sigma_{\theta}(1)}\geq\ldots\geq\theta_{\sigma_{\theta}(K)}$. Then, a high-probability set $S$ can be constructed as
$S=\{\sigma_{\theta}(1),\ldots,\sigma_{\theta}(k)\}$
for some $0\leq k<K$. Since $S$ contains the indices of the $k$ highest-valued components of $\theta^{\ast}$, it is expected to cover a large portion of the total probability for $X$. This can be the case even for a small value of $k$ relative to $K$ when the components of $\theta^{\ast}$ are not evenly distributed. Hence, a sensible alternative to SRR would be to confine the randomized response to the set $S$ (unioned by a random element from $S^{c}$ to remain LDP). The expected benefit of this approach is due to (i) using less amount of randomization since $k<K$, and (ii) having an informative response since $X$ belongs $S$ with a high probability.
Based on this approach, we propose the Randomly Restricted Randomized Response (RRRR) mechanism, given in Algorithm 2. RRRR has three algorithmic parameters: a subset $S$ of $\{1,\ldots,K\}$ and two privacy parameters $\epsilon_{1}$ and $\epsilon_{2}$ which operates on $S$ and $S^{c}$, respectively. Note that when $|S|=0$ and $\epsilon_{2}=\epsilon$, RRRR reduces to SRR. Theorem 1 states the necessary conditions for $\epsilon_{1},\epsilon$ for RRRR to be $\epsilon$-LDP. A proof of Theorem 1 is given in Appendix A.1.
Input: Sample space size $K$, a subset $S\subset\{1,\ldots,K\}$, randomization parameters $\epsilon_{1},\epsilon_{2}>0$, input $X\in\{1,\ldots,K\}$
Output: Randomized response $Y\in\{1,\ldots,K\}$
1
2Step 1: Generate $U\sim\textup{Uniform}(S^{c})$ and set
$R=U\cdot\mathbb{I}(X\in S)+\texttt{SRR}(X|\epsilon_{2},S^{c})\cdot\mathbb{I}(X\notin S).$
3Step 2: Set $Z=X\cdot\mathbb{I}(X\in S)+R\cdot\mathbb{I}(X\in S^{c})$.
4Step 3: Set $Y=\texttt{SRR}(Z|\epsilon_{1},S\cup\{R\})$.
return $Y$
Algorithm 2 RRRR: Randomly restricted randomized response mechanism
@@
We elaborate on the behavior of RRRR and the choice of $\epsilon_{1}$ and $\epsilon_{2}$ in the light of Theorem 1. The theorem implies that there is a degree of freedom to choose $\epsilon_{1}<\epsilon$, while $\epsilon_{2}$ in (2) is the largest possible choice given $\epsilon$ and $\epsilon_{1}$. In RRRR, the probability of an honest response, i.e., $X=Y$, given $X\in S$ is
$\mathbb{P}_{\text{RRRR}}(Y=X|X\in S)=\frac{e^{\epsilon_{1}}}{e^{\epsilon_{1}}+|S|},$
which should be contrasted to $\mathbb{P}_{\text{SRR}}(X=Y|X\in S)=\frac{e^{\epsilon}}{e^{\epsilon}+K-1}$ for SRR. Anticipating that $\{X\in S\}$ is likely, one should at least aim for a higher probability for $Y=X$ than that of SRR when $X\in S$ for RRRR to be relevant. That is, we would like to choose $\epsilon_{1}$ to satisfy $\mathbb{P}_{\text{RRRR}}(X=Y|X\in S)\geq\mathbb{P}_{\text{SRR}}(X=Y|X\in S)$, which is equivalent to
@@
Taking into account also the constraint that $\epsilon_{1}\leq\epsilon$ (by Theorem 1), we consider $\epsilon_{1}=\kappa\epsilon$, where $\kappa\in[0,1]$. It is possible to choose $\kappa=1$; however, again by Theorem 1, this requires that $\epsilon_{2}=0$, which renders $Y$ completely uninformative when $X\notin S$. We suggest taking $\kappa$ a number close to $1$, such as $0.9$, to ensure (3) with a significant margin. In Section 7, we discuss the choice of $\kappa$ in more detail.
Example 1 (Numerical illustration).
We illustrate the benefit of RRRR and the role of $k$ with an example. We considered $\theta$ values such that $\theta_{i}/\theta_{i+1}$ is constant for $i=1,\ldots,K-1$. The ratio $\theta_{i}/\theta_{i+1}$ controls the degree of ‘evenness’ in $\theta$: The smaller ratio indicates a more evenly distributed $\theta$. Note that $\theta$ is already ordered in this example; hence, we consider using $S=\{1,\ldots,k\}$ which has the $k$ most likely values for $X$ according to $\theta$. Also, for a given $\epsilon$, we fix $\epsilon_{1}=0.9\epsilon$ and set $\epsilon_{2}$ according to (2).
Figure 2 shows, for a fixed $\epsilon$ and $K=20$, and various values of $k$, the probability of the randomized response being equal to the sensitive information, i.e., $\mathbb{P}_{\textup{RRRR}}(Y=X|\theta)$, vs $\theta_{i}/\theta_{i+1}$ when $S=\{1,\ldots,k\}$ in RRRR. (This probability is considered one of the utility functions presented in Sections 4.2.1-4.2.2.) Comparing this probability with $\mathbb{P}_{\textup{SRR}}(Y=X|\theta)$, it can be observed that RRRR can do significantly better than SRR if $k$ can be chosen suitably. The plots demonstrate that the “suitable” $k$ depends on $\theta$: While the best $k$ tends to be larger for more even $\theta$, small $k$ becomes the better choice for non-even $\theta$ (large $\theta_{i}/\theta_{i+1}$). This is because, when $\theta_{i}/\theta_{i+1}$ is large, the probability is concentrated on just a few components, and $S$ with a small $k$ captures most of the probability. Moreover, the plots for $\epsilon=1$ and $\epsilon=5$ also show the effect of the level of privacy. In more challenging scenarios where $\epsilon$ is smaller, the gain obtained by RRRR compared to SRR is bigger.
Refer to caption Refer to caption
Figure 2: $\mathbb{P}_{\textup{RRRR}}(Y=X|\theta)$ vs $\theta_{i}/\theta_{i+1}$ for all $i=1,\ldots,K-1$ with $K=20$. Left: $\epsilon=1$, Right: $\epsilon=5$.
@@ 4.2 Subset selection for RRRR
RRRR in Algorithm 2 can be used for any $S$ by choosing $\epsilon_{1},\epsilon_{2}$ according to $|S|$, as suggested by Theorem 1. In this paper, we assume a fixed $\kappa\in(0,1)$ and set $\epsilon_{1}=\kappa\epsilon$; and we focus on the selection of $S$. (Note that one can also seek to optimize $\epsilon_{1}$, too. However, this comes with the expense of additional computation due to the requirement to explore a larger hyperparameter space.)
To formalize the discussion, let $\texttt{RRRR}(S,\epsilon)$ be the RRRR mechanism that achieves $\epsilon$-LDP by using the subset $S$ $\epsilon_{1}=\kappa\epsilon$ and let $\texttt{RRRR}(x|S,\epsilon)$ be its random output when the input is $X=x$. Furthermore, let $U(\theta,S,\epsilon)$ be the ‘utility’ (concerning inference) when the true parameter is $\theta^{\ast}=\theta$ and $Y=\texttt{RRRR}(X|S,\epsilon)$. One would like to choose $S$ such that randomized response $Y$ maximizes $U(\theta,S,\epsilon)$. However, since there are $12^{K}-$ 1 feasible choices for $S$, one must confine the search space for $S$ in practice. As suggested in Example 1, RRRR becomes most relevant when the set $S$ is chosen to have the indices of the $k$ largest components of $\theta$. Therefore, for a given $\theta$, we confine the choices for $S$ to
@@
Note that the set of solutions considered in (5) covers SRR, obtained with $k=0$ (leading to $S=\emptyset$ and $\epsilon_{2}=\epsilon$).
In the subsequent sections, we present six different utility functions and justify their relevance to estimation; the usefulness of the proposed functions is also demonstrated in the numerical experiments.
@@ 4.2.1 Fisher information matrix
The first utility function under consideration is based on the Fisher information matrix at $\theta$ according to the distribution of $Y$ given $\theta$. It is well-known that the inverse of the Fisher information matrix sets the Cramer-Rao lower bound for the variance of an unbiased estimator (Casella and Berger, , 2021). Hence the Fisher information can be regarded as a reasonable metric to quantify the information contained in $Y$ about $\theta$. This approach is adopted in Lopuhaä-Zwakenberg et al., (2022); Steinberger, (2024) for LDP applications for estimating discrete distributions, and Alparslan and Yıldırım, (2022); Yıldırım, (2024) in similar problems involving parametric continuous distributions.
For a given $\theta\in\Delta$, let $F(\theta;S,\epsilon)$ be the Fisher information matrix evaluated at $\theta$ when $X\sim\text{Cat}(\theta)$ and $Y=\texttt{RRRR}(X|S,\epsilon)$. Let $g_{S,\epsilon}$ be the conditional probability mass function of the distribution of $Y$ given $X$ according to $Y=\texttt{RRRR}(X|S,\epsilon)$, that is,
$g_{S,\epsilon}(y|x):=\mathbb{P}_{\textup{RRRR}}(Y=y|X=x)$
when $Y=\texttt{RRRR}(X|S,\epsilon)$. The following result states $F(\theta;S,\epsilon)$ in terms of $g_{S,\epsilon}$ and $\theta$. The result is derived in Lopuhaä-Zwakenberg et al., (2022); we also give a simple proof in Appendix A.2. Note that $F(\theta;S,\epsilon)$ is $(K-1)\times(K-1)$ since $\theta$ has $K-1$ free components and $\theta_{K}=1-\sum_{i=1}^{K-1}\theta_{i}$.
@@
This utility function depends on the Fisher information differently from Lopuhaä-Zwakenberg et al., (2022); Steinberger, (2024), who considered the determinant of the FIM as the utility function. The rationale behind (7) is that the for an unbiased estimator $\hat{\theta}(Y)$ of $\theta^{\ast}$ based on $Y=\texttt{RRRR}(X|S,\epsilon)$ and $X\sim\text{Cat}(\theta^{\ast})$, the expected mean squared error is bounded by $E_{\theta^{\ast}}[\|\hat{\theta}(Y)-\theta^{\ast}\|^{2}]\leq\textup{Tr}\left[F^{-1}(\theta^{\ast};S,\epsilon)\right]$. For the utility function in (7) to be well defined, the FIM needs to be invertible. Proposition 2, proven in Appendix A.2, states that this is indeed the case.
@@
$F(\theta;S,\epsilon)$ in (6) is invertible for all $\theta\in\Delta$, $S\subset\{1,\ldots,K\}$, and $\epsilon_{1},\epsilon_{2}>0$.

### v2

@@ 3 Problem definition and general framework
Suppose we are interested in a discrete probability distribution $\mathcal{P}$ of a certain form of sensitive categorical information $X\in[K]:=\{1,\ldots,K\}$ of individuals in a population. Hence, $\mathcal{P}$ is a categorical distribution $\text{Cat}(\theta^{\ast})$ with a probability vector
@@
$\Delta:=\left\{\theta\in\mathbb{R}^{K}:\sum_{k=1}^{K}\theta_{k}=1\text{ and }\theta_{k}\geq 0\text{ for }k\in[K]\right\}.$
We assume a setting where individuals’ sensitive data are collected privately and sequentially in time. The privatization is performed via a randomized algorithm that, upon taking a category index in $[K]$ as an input, returns a random category index in $[K]$ such that the whole data collection process is $\epsilon$-LDP. (See Definition 1.) Let $X_{t}$ and $Y_{t}$ be the private information and randomized responses of individual $t$, respectively. According to Definition 1 for LDP, the following inequality must be satisfied for all triples $(x,x^{\prime},y)\in[K]^{3}$ for the randomized mechanism to be $\epsilon$-LDP.
@@
This paper concerns the Bayesian estimation of $\theta$ while adapting the randomized mechanism to improve the estimation utility continually. We propose a general framework called AdOBEst-LDP, in which the randomized mechanism at time $t$ is adapted to the data collected until time $t-1$. AdOBEst-LDP is outlined in Algorithm 1.
@@ 5      Step 3: Draw a sample (approximately) from the posterior distribution
Algorithm 1 is fairly general, and it does not describe how to choose the $\epsilon$-LDP mechanism $\mathcal{M}_{t}$ at time $t$, nor does it provide the details of the posterior sampling. However, it is still worth making some critical observations about the nature of the algorithm. Firstly, at time $t$ the selection of the $\epsilon$-LDP mechanism in Step 1 relies on the posterior sample $\Theta_{t-1}$, which serves as an estimator of the true parameter $\theta^{\ast}$ based on the past observations. As we shall see in Section 4, at Step 1 the ‘best’ $\epsilon$-LDP mechanism is chosen from a set of candidate LDP mechanisms according to a utility function. This step is relevant only when $\Theta_{t-1}$ is a reliable estimator of $\theta^{\ast}$. In other words, Step 1 ‘exploits’ the estimator $\Theta_{t-1}$. Moreover, the random nature of posterior sampling prevents having too much confidence in the current estimator $\Theta_{t-1}$ and enables a certain degree of ‘exploration.’ In conclusion, Algorithm 1 utilizes an ‘exploration-exploitation’ approach reminiscent of reinforcement learning. In particular, posterior sampling in Step 3 suggests a strong parallelism between AdOBEst-LDP and the well-known exploration-exploitation approach called Thompson sampling (Russo et al., , 2018).
@@ 4 Constructing informative randomized response mechanisms
In this section, we describe Steps 1-2 of AdOBEst-LDP in Algorithm 1 where the $\epsilon$-LDP mechanism $\mathcal{M}_{t}$ is selected at time $t$ based on the posterior sample $\Theta_{t}$ and a randomized response is generated using $\mathcal{M}_{t}$. For ease of exposition, we will drop the time index $t$ throughout the section and let $\Theta_{t-1}=\theta$.
Recall from Definition 1 that an $\epsilon$-LDP randomized mechanism is associated with a conditional probability distribution that satisfies (1). An $\epsilon$-LDP mechanism is not unique. One such mechanism is the standard randomized response mechanism (SRR). For subsequent use, it is convenient to define SRR generally: We let $\texttt{SRR}(X;\Omega,\epsilon)$ the output of SRR which operates on the set $\Omega$ with LDP parameter $\epsilon$ when the input is $X\in\Omega$. Then, we have
$Y=\text{SRR}(X;\Omega,\epsilon)=\begin{cases}X&\text{w.p. }e^{\epsilon}/(e^{\epsilon}+|\Omega|-1)\\ \sim\text{Uniform}(\Omega/\{X\})&\text{ else }\end{cases}.$ (2)
We aim to develop an alternative randomized mechanism whose response $Y$ is more informative about $\theta^{\ast}$ than the one generated as $Y=\text{SRR}(X;[K],\epsilon)$. The main idea is as follows. Supposing that the posterior sample $\Theta_{t-1}=\theta$ is an accurate estimate of $\theta^{\ast}$, it is reasonable to aim for the ‘best’ $\epsilon$-LDP mechanism (among a range of candidates) which would maximize the (estimation) utility of $Y$ if the true parameter were $\theta^{\ast}=\theta$. We follow this main idea to develop the proposed $\epsilon$-LDP mechanism.
@@ 4.1 The randomly restricted randomized response (RRRR) mechanism
Given $\Theta_{t-1}=\theta\in\Delta$, an informative randomized response mechanism can be constructed by considering a high-probability set $S\subset[K]$ and a low-probability set $S^{c}=[K]/S$ for $X$ (according to $\theta$). Then, a sensible alternative to $\text{SRR}(X;[K],\epsilon)$ would be to confine the randomized response to the set $S$ (unioned by a random element from $S^{c}$ to remain LDP). The expected benefit of this approach is due to (i) using less amount of randomization since $|S|<K$, and thus (ii) having an informative response when $X\in S$, which happens with a high probability. Based on this approach, we propose RRRR, whose precise steps are given in Algorithm 2.
Input: Sample space size $K$, a subset $S\subset[K]$, privacy parameters $\epsilon_{1},\epsilon_{2}>0$, input $X\in[K]$
Output: Randomized response $Y\in[K]$
1 if $X\in S$ then
2       Draw $R\sim\textup{Uniform}(S^{c})$.
3      Set $Y=\texttt{SRR}(X;S\cup\{R\},\epsilon_{1})$ as in (2).
4 else
5       Set $R=\texttt{SRR}(X;S^{c},\epsilon_{2})$ as in (2).
6      Set $Y=\texttt{SRR}(R;S\cup\{R\},\epsilon_{1})$ as in (2).
7return $Y$
Algorithm 2 RRRR
RRRR has three algorithmic parameters: a subset $S$ of $[K]$ and two privacy parameters $\epsilon_{1}$ and $\epsilon_{2}$ which operates on $S$ and $S^{c}$, respectively. Theorem 1 states the necessary conditions for $\epsilon_{1}$ and $\epsilon_{2}$ for RRRR to be $\epsilon$-LDP. A proof of Theorem 1 is given in Appendix A.1.
@@
Note that when $S=\emptyset$ and $\epsilon_{2}=\epsilon$, RRRR reduces to SRR.
4.2 Choosing the privacy parameters $\epsilon_{1}$, $\epsilon_{2}$
We elaborate on the choice of $\epsilon_{1}$ and $\epsilon_{2}$ in the light of Theorem 1. In RRRR, the probability of an honest response, i.e., $X=Y$, given $X\in S$, is
$\mathbb{P}(Y=X|X\in S)=\frac{e^{\epsilon_{1}}}{e^{\epsilon_{1}}+|S|},$
which should be contrasted to $e^{\epsilon}/(e^{\epsilon}+K-1)$, which would be the probability if $Y=\text{SRR}(X;[K],\epsilon)$. Anticipating that $\{X\in S\}$ is likely, one should at least aim for $\epsilon_{1}$ that satisfies $\mathbb{P}(X=Y|X\in S)\geq e^{\epsilon}/(e^{\epsilon}+K-1)$ for RRRR to be relevant. This is equivalent to
@@ 4.2 Choosing the privacy parameters $\epsilon_{1}$, $\epsilon_{2}$
Taking into account also the constraint that $\epsilon_{1}\leq\epsilon$ (by Theorem 1), we suggest $\epsilon_{1}=\kappa\epsilon$, where $\kappa\in(0,1)$ is a number close to $1$, such as $0.9$, to ensure (4) with a significant margin. (It is possible to choose $\kappa=1$; however, again by Theorem 1, this requires that $\epsilon_{2}=0$, which renders $Y$ completely uninformative when $X\notin S$.) In Section 7, we discuss the choice of $\kappa$ in more detail.
For the next section, we assume a fixed $\kappa\in(0,1)$, and set $\epsilon_{1}=\kappa\epsilon$; and we focus on the selection of $S$.
@@ 4.3 Subset selection for RRRR
Let $\texttt{RRRR}(X;S,\epsilon)$ be the random output of RRRR that achieves $\epsilon$-LDP by using the subset $S$ and the privacy parameters $\epsilon_{1}=\kappa\epsilon$ and $\epsilon_{2}$ as in (3) when the input is $X$. Furthermore, let $U(\theta,S,\epsilon)$ be the (inferential) ‘utility’ of $Y=\texttt{RRRR}(X;S,\epsilon)$ when $X\sim\text{Cat}(\theta)$. One would like to choose $S$ that maximizes $U(\theta,S,\epsilon)$. (One could also seek to optimize $\kappa$ in $\epsilon_{1}=\kappa\epsilon$, too, however with the expense of additional computation.)
However, since there are $12^{K}-$ 1 feasible choices for $S$, one must confine the search space for $S$ in practice. As discussed above, RRRR becomes most relevant when the set $S$ is a high-probability set. Therefore, for a given $\theta$, we confine the choices for $S$ to
@@
where $\sigma_{\theta}:=(\sigma_{\theta}(1),\ldots,\sigma_{\theta}(K))$ be the permutation vector for $\theta$ so that $\theta_{\sigma_{\theta}(1)}\geq\ldots\geq\theta_{\sigma_{\theta}(K)}$.
@@
The alternatives in (5) can be justified. Since $S_{\theta,k}$ contains the indices of the $k$ highest-valued components of $\theta^{\ast}$, it is expected to cover a large portion of the total probability for $X$. This can be the case even for a small value of $k$ relative to $K$ when the components of $\theta^{\ast}$ are not evenly distributed. Also, the alternatives cover the basic SRR, which is obtained with $k=0$ (leading to $S=\emptyset$ and $\epsilon_{2}=\epsilon$).
In the subsequent sections, we present six different utility functions $U(\theta,S,\epsilon)$ and justify their relevance to estimation; the usefulness of the proposed functions is also demonstrated in the numerical experiments.
@@ 4.3.1 Fisher information matrix
The first utility function under consideration is based on the Fisher information matrix at $\theta$ according to the distribution of $Y$ given $\theta$. It is well-known that the inverse of the Fisher information matrix sets the Cramer-Rao lower bound for the variance of an unbiased estimator. Hence, the Fisher information can be regarded as a reasonable metric to quantify the information contained in $Y$ about $\theta$. This approach is adopted in Lopuhaä-Zwakenberg et al., (2022); Steinberger, (2024) for LDP applications for estimating discrete distributions, and Alparslan and Yıldırım, (2022); Yıldırım, (2024) in similar problems involving parametric continuous distributions.
For a given $\theta\in\Delta$, let $F(\theta;S,\epsilon)$ be the Fisher information matrix evaluated at $\theta$ when $X\sim\text{Cat}(\theta)$ and $Y=\texttt{RRRR}(X;S,\epsilon)$. Let
$g_{S,\epsilon}(y|x):=\mathbb{P}(Y=y|X=x)$
when $Y=\texttt{RRRR}(X;S,\epsilon)$. The following result states $F(\theta;S,\epsilon)$ in terms of $g_{S,\epsilon}$ and $\theta$. The result is derived in Lopuhaä-Zwakenberg et al., (2022); we also give a simple proof in Appendix A.2. Note that $F(\theta;S,\epsilon)$ is $(K-1)\times(K-1)$ since $\theta$ has $K-1$ free components and $\theta_{K}=1-\sum_{i=1}^{K-1}\theta_{i}$.
@@
This utility function depends on the Fisher information differently from Lopuhaä-Zwakenberg et al., (2022); Steinberger, (2024), who considered the determinant of the FIM as the utility function. The rationale behind (8) is that the for an unbiased estimator $\hat{\theta}(Y)$ of $\theta^{\ast}$ based on $Y=\texttt{RRRR}(X;S,\epsilon)$, the expected mean squared error is bounded by $E_{\theta^{\ast}}[\|\hat{\theta}(Y)-\theta^{\ast}\|^{2}]\leq\textup{Tr}\left[F^{-1}(\theta^{\ast};S,\epsilon)\right]$. For the utility function in (8) to be well-defined, the FIM needs to be invertible. Proposition 2, proven in Appendix A.2, states that this is indeed the case.
@@
$F(\theta;S,\epsilon)$ in (7) is invertible for all $\theta\in\Delta$, $S\subset[K]$, and $\epsilon_{1},\epsilon_{2}>0$.